# On an unconditional $\rm GL_{3}$ analog of Selberg’s result

### v1

@@ 3. An approximation of $S_{F}(t)$
With the above observation and the Hecke relations, the numerator of this general term (4.4) can be written as a sum whose terms are all of the form $A_{F}(m_{1},n_{1})\overline{A_{F}(m_{2},n_{2})}$ where $(m_{1},n_{1})\neq(m_{2},n_{2})$. Since for these terms $\delta_{n_{1},m_{1}}\delta_{n_{2},m_{2}}=0$, we see from Theorem 2.2 and the estimate (1.8) that the contribution from the general term (4.4) to the moment (4.2) is
$\begin{split}&\ll_{n}\frac{1}{\mathcal{H}}\sum_{p_{1}<p_{2}<\cdots<p_{r}\leq T^{\delta}}p_{1}^{-\frac{1}{2}(m(p_{1})+n(p_{1}))}\cdots p_{r}^{-\frac{1}{2}(m(p_{r})+n(p_{r}))}\\ &\qquad\qquad\qquad\qquad\qquad\times\left((T^{1+2n\delta})^{\varepsilon}\big{(}T^{1+n\delta}+T^{3}+T^{1+2n\delta\vartheta}M^{2}\big{)}\right)\\ &\ll_{n}\frac{1}{\mathcal{H}}\bigg{(}\sum_{p\leq T^{\delta}}p^{-\frac{1}{2}}\bigg{)}^{r}\big{(}T^{1+\varepsilon+n\delta(1+2\varepsilon)}+T^{3+\varepsilon+2n\delta\varepsilon}+T^{1+\varepsilon+2n\delta(\vartheta+\varepsilon)}M^{2}\big{)}\\ &\ll_{n}T^{-2+\varepsilon+n\delta(3/2+2\varepsilon)}M^{-2}+T^{\varepsilon+n\delta(1/2+2\varepsilon)}M^{-2}+T^{-2+\varepsilon+n\delta(1/2+2\vartheta+2\varepsilon)},\end{split}$
@@
[…]
@@
$\begin{split}&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\binom{2m}{m}(m!)(m!)\frac{(-1)^{m}(-i)^{2m}}{(2\pi)^{2m}}\\ &\qquad\sum_{p_{1}<p_{2}<\cdots<p_{m}\leq T^{\delta}}\frac{A_{F}(1,p_{1})\overline{A_{F}(1,p_{1})}\cdots A_{F}(1,p_{m})\overline{A_{F}(1,p_{m})}}{p_{1}\cdots p_{m}}.\end{split}$ (4.6)
@@
Applying Theorem 2.2 and the estimate (4.5), we have (4.6) equals
$\begin{split}&\frac{1}{\mathcal{H}}\frac{(2m)!}{(2\pi)^{2m}}\sum_{p_{1}<p_{2}<\cdots<p_{m}\leq T^{\delta}}\frac{1}{p_{1}\cdots p_{m}}\cdot\left(\mathcal{H}+O\left(T^{1+\varepsilon}(p_{1}\cdots p_{m})+T^{3+\varepsilon}+T^{1+\varepsilon}M^{2}(p_{1}\cdots p_{m})^{2\vartheta}\right)\right)\\ =&\frac{(2m)!}{m!(2\pi)^{2m}}\sum_{p_{1},p_{2},\ldots,p_{m}\leq T^{\delta}\atop p_{j}\,\text{distinct}}\frac{1}{p_{1}\cdots p_{m}}+O\bigg{(}T^{-2+\varepsilon}M^{-2}\sum_{p_{1},p_{2},\ldots,p_{m}\leq T^{\delta}\atop p_{j}\,\text{distinct}}1\bigg{)}\\ &+O\bigg{(}T^{\varepsilon}M^{-2}\sum_{p_{1},p_{2},\ldots,p_{m}\leq T^{\delta}\atop p_{j}\,\text{distinct}}\frac{1}{p_{1}\cdots p_{m}}\bigg{)}+O\bigg{(}T^{-2+\varepsilon}\sum_{p_{1},p_{2},\ldots,p_{m}\leq T^{\delta}\atop p_{j}\,\text{distinct}}\frac{1}{(p_{1}\cdots p_{m})^{1-2\vartheta}}\bigg{)}\\ \\ =&\frac{(2m)!}{m!(2\pi)^{2m}}(\log\log T+O(1))^{m}+O\big{(}T^{m\delta-2+\varepsilon}M^{-2}+T^{\varepsilon}M^{-2}(\log\log T)^{m}+T^{2m\delta\vartheta-2+\varepsilon}(\log\log T)^{m}\big{)}\\ =&\frac{(2m)!}{m!(2\pi)^{2m}}(\log\log T)^{m}+O_{m}\big{(}(\log\log T)^{m-1}\big{)}.\end{split}$
@@
$\begin{split}&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}|\mathcal{R}_{F}(t)|^{2n}\\ \ll&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\left|\operatorname{Im}\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+it}\log p}\right|^{2n}+\sum_{k=2}^{7}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\left|\operatorname{Im}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+it)}\log p}\right|^{2n}\\ &+\sum_{k=1}^{7}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}(\sigma_{x}-1/2)^{2n}x^{2nk(\sigma_{x}-1/2)}\left(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|\mathrm{d}\sigma\right)^{2n}\\ &+\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}(\sigma_{x}-1/2)^{2n}\big{(}\log(|t|+M_{F})\big{)}^{2n}+\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}.\end{split}$ (4.7)
For the first two terms in (4.7), we have
$\begin{split}&\left|\operatorname{Im}\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+it}\log p}\right|^{2n}\ll_{n}\left(\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+it}\log p}-\sum_{p\leq x^{3}}\frac{\overline{A_{F}(1,p)}(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2-it}\log p}\right)^{2n},\end{split}$
@@
$\begin{split}\left|\operatorname{Im}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+it)}\log p}\right|^{2n}\ll_{n}\left(\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+it)}\log p}-\sum_{p^{k}\leq x^{3}}\frac{\overline{C_{F}(p^{k})}\Lambda_{x}(p^{k})}{p^{k(1/2-it)}\log p}\right)^{2n}.\end{split}$ (4.8)
@@
is a symmetric polynomial in $\alpha_{F,j}(p)$, since (1.5) says that $1$, $A_{F}(1,p)$, $A_{F}(p,1)=\overline{A_{F}(1,p)}$ are the elementary symmetric polynomials in $\alpha_{F,j}(p)$, the fundamental theorem of symmetric polynomials (see [15, pg. 21, Remark]) implies that the symmetric polynomial $C_{F}(p^{k})=\alpha_{F,1}(p)^{k}+\alpha_{F,2}(p)^{k}+\alpha_{F,3}(p)^{k}$ is a polynomial in $1$, $A_{F}(1,p)$ and $\overline{A_{F}(1,p)}$. Then we can use the fact $|\Lambda_{x}(p)-\Lambda(p)|=O\left(\frac{\log^{3}p}{\log^{2}x}\right)$ and the similar argument as Case (I)–(III) to see that the contributions from first two terms in (4.7) to (4.3) is of $O_{t,n}(1)$. The contributions from last two terms in (4.7) is of $O_{t,n}(1)$ by Cauchy–Schwarz inequality, the assumption (1.7), Corollary 2.3 and Lemma 4.1. For the third term in (4.7), it follows from Cauchy–Schwarz inequality that
$\begin{split}&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}(\sigma_{x}-1/2)^{2n}x^{2nk(\sigma_{x}-1/2)}\left(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|\mathrm{d}\sigma\right)^{2n}\\ &\leq\left(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}(\sigma_{x}-1/2)^{4n}x^{4nk(\sigma_{x}-1/2)}\right)^{1/2}\\ &\times\left(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\left(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|\mathrm{d}\sigma\right)^{4n}\right)^{1/2},\end{split}$ (4.9)
where $1\leq k\leq 7$. By Hölder’s inequality with the exponents $4n/(4n-1)$ and $4n$,
$\begin{split}&\left(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|\mathrm{d}\sigma\right)^{4n}\\ \leq&\left(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\mathrm{d}\sigma\right)^{4n-1}\left(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|^{4n}\mathrm{d}\sigma\right)\\ =&\frac{1}{(k\log x)^{4n-1}}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|^{4n}\mathrm{d}\sigma.\end{split}$
@@
$\begin{split}&\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|^{4n}\\ &\ll_{n}\left(\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}+\sum_{p^{k}\leq x^{3}}\frac{\overline{C_{F}(p^{k})}\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma-it)}}\right)^{4n}\\ &\quad+\left(\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}-\sum_{p^{k}\leq x^{3}}\frac{\overline{C_{F}(p^{k})}\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma-it)}}\right)^{4n}.\end{split}$
@@
$\begin{split}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|^{4n}&\ll_{t,n}(\log x)^{8n}.\end{split}$
@@
$\begin{split}&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\left(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\left|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+it)}}\right|\mathrm{d}\sigma\right)^{4n}\\ \ll_{t,n}\,&\frac{1}{(k\log x)^{4n-1}}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}(\log x)^{8n}\mathrm{d}\sigma\ll_{t,n}(\log x)^{4n}.\end{split}$
Inserting this bound to (4.9) and using Lemma 4.1, we have the third term in (4.7) is of $O_{t,n}(1)$. Thus we complete the proof of this lemma. ∎
@@
$\begin{split}S_{F}(t)^{n}=\mathcal{M}_{F}(t)^{n}+O_{n}\left(\sum_{\ell=1}^{n}|\mathcal{M}_{F}(t)|^{n-\ell}|\mathcal{R}_{F}(t)|^{\ell}\right),\end{split}$ (4.10)
@@
$\begin{split}&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}|\mathcal{M}_{F}(t)|^{n-\ell}|\mathcal{R}_{F}(t)|^{\ell}\\ \ll&\left(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\right)^{\frac{1}{2}}\left(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}|\mathcal{M}_{F}(t)|^{2n}\right)^{\frac{n-\ell}{2n}}\\ &\times\left(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}|\mathcal{R}_{F}(t)|^{2n}\right)^{\frac{\ell}{2n}}\\ \ll&{}_{t,n}\,(\log\log T)^{(n-\ell)/2}\ll_{t,n}(\log\log T)^{(n-1)/2}.\end{split}$
Here we use the facts that the positivity of $\mathcal{N}_{F}$ and $h_{T,M}(\mu_{F})\geq 0$ for all $\rm GL_{3}$ Hecke–Maass cusp forms $F$. Then (1.10) follows from (4.2), (4.10) and the above estimate.
@@
$\int_{\mathbb{R}}\xi^{n}\mathrm{d}\mu_{TM}(\xi)=\left(\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\left(\frac{S_{F}(t)}{\sqrt{\log\log T}}\right)^{n}\right)\bigg{/}\left(\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\right).$

### v2

@@ 3. An approximation of $S_{F}(t)$
With the above observation and the Hecke relations, the numerator of this general term (4.4) can be written as a sum whose terms are all of the form $A_{F}(m_{1},n_{1})\overline{A_{F}(m_{2},n_{2})}$ where $(m_{1},n_{1})\neq(m_{2},n_{2})$. Since for these terms $\delta_{m_{1},m_{2}}\delta_{n_{1},n_{2}}=0$, we see from Theorem 2.2 and the estimate (1.9) that the contribution from the general term (4.4) to the moment (4.2) is
$\begin{split}&\ll_{n}\frac{1}{\mathcal{H}}\sum_{p_{1}<p_{2}<\cdots<p_{r}\leq T^{\delta}}p_{1}^{-\frac{1}{2}(m(p_{1})+n(p_{1}))}\cdots p_{r}^{-\frac{1}{2}(m(p_{r})+n(p_{r}))}\\ &\qquad\qquad\qquad\qquad\qquad\times\bigg((T^{1+2n\delta})^{\varepsilon}\big(T^{1+n\delta}+T^{3}+T^{1+2n\delta\vartheta}M^{2}\big)\bigg)\\ &\ll_{n}\frac{1}{\mathcal{H}}\bigg(\sum_{p\leq T^{\delta}}p^{-\frac{1}{2}}\bigg)^{r}\big(T^{1+\varepsilon+n\delta(1+2\varepsilon)}+T^{3+\varepsilon+2n\delta\varepsilon}+T^{1+\varepsilon+2n\delta(\vartheta+\varepsilon)}M^{2}\big)\\ &\ll_{n}T^{-2+\varepsilon+n\delta(3/2+2\varepsilon)}M^{-2}+T^{\varepsilon+n\delta(1/2+2\varepsilon)}M^{-2}+T^{-2+\varepsilon+n\delta(1/2+2\vartheta+2\varepsilon)},\end{split}$
@@
Case (II): In the general term (4.4), $m(p_{j})\equiv n(p_{j})(\bmod\,3)$ for all $j$, and $m(p_{j_{0}})\geq 2$ or $n(p_{j_{0}})\geq 2$ for some $j_{0}$. In this case, $m(p_{j})+n(p_{j})\geq 2$ for all $j$ and $m(p_{j_{0}})+n(p_{j_{0}})\geq 3$ for some $j_{0}$, so $n>2r$ and $r<\lfloor\frac{n-1}{2}\rfloor$. By applying Lemma 2.1 to the numerator of this general term (4.4) and multiplying out by Hecke relations, one sees that it is possible to have resulting terms $A_{F}(m_{1},n_{1})\overline{A_{F}(m_{2},n_{2})}$ where $(m_{1},n_{1})=(m_{2},n_{2})$. Using Theorem 2.2, the contribution to the moment (4.2) such terms is bounded by the diagonal term of size $\asymp T^{3}M^{2}$. The contribution from the off-diagonal terms is clearly negligible compared to $T^{3}M^{2}$, given that $0<\delta<\frac{1}{n}$. Using the classical result
@@
$\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\binom{2m}{m}(m!)(m!)\frac{(-1)^{m}(-\mathrm{i})^{2m}}{(2\pi)^{2m}}$
$\qquad\sum_{p_{1}<p_{2}<\cdots<p_{m}\leq T^{\delta}}\frac{A_{F}(1,p_{1})\overline{A_{F}(1,p_{1})}\cdots A_{F}(1,p_{m})\overline{A_{F}(1,p_{m})}}{p_{1}\cdots p_{m}}.$ (4.6)
@@
Applying Theorem 2.2 and the estimate (4.5), we have (4) equals
$\begin{split}&\frac{1}{\mathcal{H}}\frac{(2m)!}{(2\pi)^{2m}}\sum_{p_{1}<p_{2}<\cdots<p_{m}\leq T^{\delta}}\frac{1}{p_{1}\cdots p_{m}}\cdot\bigg(\mathcal{H}+O\bigg(T^{1+\varepsilon}(p_{1}\cdots p_{m})+T^{3+\varepsilon}+T^{1+\varepsilon}M^{2}(p_{1}\cdots p_{m})^{2\vartheta}\bigg)\bigg)\\ =&\frac{(2m)!}{m!(2\pi)^{2m}}\sum_{p_{1},p_{2},\ldots,p_{m}\leq T^{\delta}\atop p_{j}\,\text{distinct}}\frac{1}{p_{1}\cdots p_{m}}+O\bigg(T^{-2+\varepsilon}M^{-2}\sum_{p_{1},p_{2},\ldots,p_{m}\leq T^{\delta}\atop p_{j}\,\text{distinct}}1\bigg)\\ &+O\bigg(T^{\varepsilon}M^{-2}\sum_{p_{1},p_{2},\ldots,p_{m}\leq T^{\delta}\atop p_{j}\,\text{distinct}}\frac{1}{p_{1}\cdots p_{m}}\bigg)+O\bigg(T^{-2+\varepsilon}\sum_{p_{1},p_{2},\ldots,p_{m}\leq T^{\delta}\atop p_{j}\,\text{distinct}}\frac{1}{(p_{1}\cdots p_{m})^{1-2\vartheta}}\bigg)\\ =&\frac{(2m)!}{m!(2\pi)^{2m}}(\log\log T+O(1))^{m}+O\big(T^{m\delta-2+\varepsilon}M^{-2}+T^{\varepsilon}M^{-2}(\log\log T)^{m}+T^{2m\delta\vartheta-2+\varepsilon}(\log\log T)^{m}\big)\\ =&\frac{(2m)!}{m!(2\pi)^{2m}}(\log\log T)^{m}+O_{m}\big((\log\log T)^{m-1}\big).\end{split}$
@@
$\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}|\mathcal{R}_{F}(t)|^{2n}$
$\ll$ $\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg|\operatorname{Im}\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+\mathrm{i}t}\log p}\bigg|^{2n}+\sum_{k=2}^{7}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg|\operatorname{Im}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+\mathrm{i}t)}\log p}\bigg|^{2n}$
$+$ $\sum_{k=1}^{7}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}(\sigma_{x}-1/2)^{2n}x^{2nk(\sigma_{x}-1/2)}\bigg(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|\mathrm{d}\sigma\bigg)^{2n}$
$+$ $\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}(\sigma_{x}-1/2)^{2n}\big(\log(|t|+M_{F})\big)^{2n}+\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}.$ (4.7)
For the first two terms in (4), we have
$\begin{split}&\bigg|\operatorname{Im}\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+\mathrm{i}t}\log p}\bigg|^{2n}\\ &\qquad\ll_{n}\bigg(\sum_{p\leq x^{3}}\frac{A_{F}(1,p)(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2+\mathrm{i}t}\log p}-\sum_{p\leq x^{3}}\frac{\overline{A_{F}(1,p)}(\Lambda_{x}(p)-\Lambda(p))}{p^{1/2-\mathrm{i}t}\log p}\bigg)^{2n},\end{split}$
@@
$\begin{split}\bigg|\operatorname{Im}\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+\mathrm{i}t)}\log p}\bigg|^{2n}\ll_{n}\bigg(\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})}{p^{k(1/2+\mathrm{i}t)}\log p}-\sum_{p^{k}\leq x^{3}}\frac{\overline{C_{F}(p^{k})}\Lambda_{x}(p^{k})}{p^{k(1/2-\mathrm{i}t)}\log p}\bigg)^{2n}.\end{split}$ (4.8)
@@
is a symmetric polynomial in $\alpha_{F,j}(p)$, since (1.5) says that $1$, $A_{F}(1,p)$, $A_{F}(p,1)=\overline{A_{F}(1,p)}$ are the elementary symmetric polynomials in $\alpha_{F,j}(p)$, the fundamental theorem of symmetric polynomials (see [17, pg. 21, Remark]) implies that the symmetric polynomial $C_{F}(p^{k})=\alpha_{F,1}(p)^{k}+\alpha_{F,2}(p)^{k}+\alpha_{F,3}(p)^{k}$ is a polynomial in $1$, $A_{F}(1,p)$ and $\overline{A_{F}(1,p)}$. Then we can use the fact $|\Lambda_{x}(p)-\Lambda(p)|=O\big(\frac{\log^{3}p}{\log^{2}x}\big)$ and the similar argument as Case (I)–(III) to see that the contributions from first two terms in (4) to (4.3) are of $O_{t,n}(1)$. The contributions from last two terms in (4) are of $O_{t,n}(1)$ by Cauchy–Schwarz inequality, the assumption (1.8), Corollary 2.3 and Lemma 4.1. For the third term in (4), it follows from Cauchy–Schwarz inequality that
$\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}(\sigma_{x}-1/2)^{2n}x^{2nk(\sigma_{x}-1/2)}\bigg(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|\mathrm{d}\sigma\bigg)^{2n}$
$\leq\bigg(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}(\sigma_{x}-1/2)^{4n}x^{4nk(\sigma_{x}-1/2)}\bigg)^{1/2}$
$\times\bigg(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|\mathrm{d}\sigma\bigg)^{4n}\bigg)^{1/2},$ (4.9)
where $1\leq k\leq 7$. By Hölder’s inequality with the exponents 4n/(4n1)4n/(4n-1) and $4n$,
$\begin{split}&\bigg(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|\mathrm{d}\sigma\bigg)^{4n}\\ \leq&\,(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\mathrm{d}\sigma\bigg)^{4n-1}\bigg(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|^{4n}\mathrm{d}\sigma\bigg)\\ =&\,\frac{1}{(k\log x)^{4n-1}}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|^{4n}\mathrm{d}\sigma.\end{split}$
@@
$\begin{split}&\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|^{4n}\\ &\ll_{n}\bigg(\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}+\sum_{p^{k}\leq x^{3}}\frac{\overline{C_{F}(p^{k})}\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma-\mathrm{i}t)}}\bigg)^{4n}\\ &\quad+\bigg(\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}-\sum_{p^{k}\leq x^{3}}\frac{\overline{C_{F}(p^{k})}\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma-\mathrm{i}t)}}\bigg)^{4n}.\end{split}$
@@
$\begin{split}\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|^{4n}&\ll_{t,n}(\log x)^{8n}.\end{split}$
@@
$\begin{split}&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg(\int_{1/2}^{\infty}x^{k(1/2-\sigma)}\bigg|\sum_{p^{k}\leq x^{3}}\frac{C_{F}(p^{k})\Lambda_{x}(p^{k})\log(xp)}{p^{k(\sigma+\mathrm{i}t)}}\bigg|\mathrm{d}\sigma\bigg)^{4n}\\ &\ll_{t,n}\,\frac{1}{(k\log x)^{4n-1}}\int_{1/2}^{\infty}x^{k(1/2-\sigma)}(\log x)^{8n}\mathrm{d}\sigma\ll_{t,n}(\log x)^{4n}.\end{split}$
Inserting this bound to (4) and using Lemma 4.1, we have the third term in (4) is of $O_{t,n}(1)$. This completes the proof of the lemma. ∎
@@
$\begin{split}S_{F}(t)^{n}=\mathcal{M}_{F}(t)^{n}+O_{n}\bigg(\sum_{\ell=1}^{n}|\mathcal{M}_{F}(t)|^{n-\ell}|\mathcal{R}_{F}(t)|^{\ell}\bigg),\end{split}$ (4.10)
@@
$\begin{split}&\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}|\mathcal{M}_{F}(t)|^{n-\ell}|\mathcal{R}_{F}(t)|^{\ell}\\ \ll&\bigg(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg)^{\frac{1}{2}}\bigg(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}|\mathcal{M}_{F}(t)|^{2n}\bigg)^{\frac{n-\ell}{2n}}\\ &\times\bigg(\frac{1}{\mathcal{H}}\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}|\mathcal{R}_{F}(t)|^{2n}\bigg)^{\frac{\ell}{2n}}\\ \ll&{}_{t,n}\,(\log\log T)^{(n-\ell)/2}\ll_{t,n}(\log\log T)^{(n-1)/2}.\end{split}$
Here we have used the positivity of $\mathcal{N}_{F}$ and the fact that $h_{T,M}(\mu_{F})\geq 0$ for all $\mathrm{GL}_{3}$ Hecke–Maass cusp forms $F$. Then (1.11) follows from (4.2), (4.10) and the above estimate.
@@
$\int_{\mathbb{R}}\xi^{n}\mathrm{d}\mu_{TM}(\xi)=\bigg(\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg(\frac{S_{F}(t)}{\sqrt{\log\log T}}\bigg)^{n}\bigg)\bigg/\bigg(\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg).$